Calculate the standard deviation for the two stocks

Consider the following information:

 



Probability of State
Rate of Return if State Occurs
Economyof EconomyStock AStock B
  Recession.24              .055         –.34         
  Normal.64              .135         .24         
  Boom.12              .230         .47         

 

a.

Calculate the expected return for the two stocks. (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)

b.Calculate the standard deviation for the two stocks. (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

Consider the following information:



Rate of Return if State Occurs
State ofProbability of State
Economyof EconomyStock AStock B
  Recession.24              .055         –.34         
  Normal.64              .135         .24         
  Boom.12              .230         .47         

Requirement 1= Expected Return of=

Stock A= (0.24*0.055) + (0.64*0.135) + (0.12*0.230) = 12.72%

Stock B=(0.24*-0.34) + (0.64*0.24) + (0.12*0.47) = 12.84%

Requirement 2=The Standard Deviation for the two Stocks:

Stock A=

ProbabilityRate of ReturnXPX2P
PX

0.240.0550.01320.000726
0.640.1350.08640.011664
0.120.2300.02760.006348


0.12720.018738

Standard Deviation = Square Root (0.018738- (0.1272)2 )

   = Square Root (0.018738-0.01617984) = Square Root (0.00255816)

= 0.0505 = 5.05 %

Stock B=

ProbabilityRate of ReturnXPX2P
PX

0.24-0.34-0.08160.027744
0.640.240.15360.036864
0.120.470.05640.026508


0.12840.091116

Standard Deviation = Square Root (0.091116- (0.1284)2 )

   = Square Root (0.091116-0.01648656) = Square Root (0.07462944)

= 0.27318 = 27.32 %


answered by: moon